OpenStudy (anonymous):

How do you solve 4(x + 6) = 6(x − 2) I need help!!!

OpenStudy (anonymous):

I believe you take \[4times x\] and \[4\times 6\] \[6timesx\] and \[6\times2\] then combine like terms

OpenStudy (anonymous):

4(x+6)=6(x-2) 4x+24=6x-12 2x=36 x=18

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

Thank you guys :D

OpenStudy (anonymous):

You're welcome :)